OpenStudy (anonymous):

I need help with this can someone help me? 1. Suppose a and b are real numbers where a > 0 and b < 0. a. Is –a + b positive or negative. Explain how you know. b. Is a – b positive or negative? Explain how you know. c. Is b – a positive or negative. Explain how you know.

OpenStudy (anonymous):

 b. \[if \left| b \right|=a,then~a-b=0\] \[if ~\left| b \right|>a,then~a-b ~is~negative.\] \[if~\left| b \right|<a,then~a-b~is~positive.\]

OpenStudy (anonymous):

c is same as a.

OpenStudy (anonymous):

ok